OpenStudy (anonymous):

Write y=(-3/4)x+3 in standard form using integers. Please help!! I'll give a medal!!

OpenStudy (anonymous):

first put x and y on the same side: y + (3/4)x=3 multiply both sides by 4: 3x + 4y = 3

OpenStudy (anonymous):

That's not one of the choices:/ But there is 3x-4y=12?

OpenStudy (triciaal):

y =-3/4 x + 3 to get rid of fraction multiply each entry by lcd 4y = -3x + 12